OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

put y=125000 x - 62,500 = 0.25(125000) x=0.25(125000)+62,500 which is the number of votes Maya received

OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (anonymous):

Put x=100, we can then get C which is the amount that Chris will pay for the coat

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you are welcome